40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Mar 15 14:25:05 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations Screen_Shot_2018-03-15_at_2.53.16_PM.png
    Reply  Thu Mar 15 15:58:02 2018, Craig, DailyProgress, BEAT, Cavity Power fluctuations vs Temp Fluctuations Beatnote_ASD_gpstime_1205189250.pdf
    Reply  Thu Mar 15 18:37:06 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations NitrogenCylinderRegulatorBefore.jpgWallAirGaugeBefore.jpgNitrogenCylinderRegulatorAfter.jpgScreen_Shot_2018-03-15_at_6.43.59_PM.png
       Reply  Thu Mar 15 22:48:21 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations 
       Reply  Mon Mar 19 12:10:27 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations AirSpringPressureRegulator_20180319.jpg
          Reply  Mon Mar 19 13:29:24 2018, awade, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations 
             Reply  Mon Mar 19 14:22:57 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations NorthAirSpringTubing2.jpgNorthAirSpringTubing.jpg
                Reply  Fri Mar 30 20:12:22 2018, awade, DailyProgress, Other, Disabling the air springs on the vac can: Shimming and clamping 2018-03-30_17.24.30.jpg2018-03-30_17.24.24.jpg
Message ID: 2140     Entry time: Thu Mar 15 22:48:21 2018     In reply to: 2138
Author: Craig 
Type: DailyProgress 
Category: Other 
Subject: Cavity Power fluctuations vs Temp Fluctuations 

Quick note on all this:  It seems to take the air springs a very long time to come to a steady state, i.e. the air springs are severely over-coupled. 

Before when I posted the plots showing the completely steady TRANS and REFL DC stuff, I had forgotten to remove the shims I had used to steady the vaccan.  When I removed them, the air springs took over fully supporting the vaccan, and began slowly drifting again, hurting the alignment into the cavities again.  This time, though, there was no fifteen minute oscillation, because our cylinder is outputing 40 psi no matter what.

So I took to making small adjustments to our regulator to get good alignment. This worked, but over a long period of time (~30 minutes) TRANS and REFL began decaying (TRANS decreasing, REFL increasing) due to poor alignment.  I'm still playing this game of making very fine air pressure adjustment, wait for a long time, then see where we're at as far as vaccan alignment.  I'll keep playing until we're at a steady state with the best alignment, then we can revert to realigning the periscopes.

Quote:

I added a pressured pipe T to the nitrogen cylinder in our lab so we could float the table and vaccan air springs from the same cylinder, as opposed to before when the air springs were floated from the wall.  This way we know that the air pressure is not changing every fifteen minutes.

The cylinder pressure is ~1800 psi right now, and I set the pressure regulator to 40 psi again (I took before and after pictures, pics 1 and 3).  I also took a picture of the wall gauge before I disconnected the air springs from it, it was set to ~40 psi as well.

The final plot is the same nine channels I looked at this morning, 25 minutes of data.  It seems our guess was right, because there are no longer ~15 minute TRANS and REFL waves.  Cool.

Quote:

Our initial impression of cavity power fluctuations was that temperature fluctuations in the EAOMs were the cause.  To check this, I made some REFL DC monitors yesterday.
Plotted is one hour of data from today.  Some notes:
1) TRANS DC and REFL DC for both cavites are breathing together every fifteen minutes, and are anticorrelated (one goes up, the other goes down). 
2) The temperature monitors are not fluctuating with the same regularity as the power monitors.
3) The REFL DC for the north PMC is fluctuating with temperature

This leads me to believe our power fluctuations are caused by changes of alignment into the cavities from the air springs holding up our vacuum can. 
Right now, the air springs are hooked up to the wall.  There is probably some pressure regulator which switches on and off every fifteen minutes.  To fix this, I'm going to switch the vacuum air springs over to our nitrogen cylinder we have in the lab and see if the fluctuations go away.

 

 

ELOG V3.1.3-